OpenStudy (anonymous):

The path of a ball in flight is p(x)=-0.23(x-3.4)^2+4.2, where x is the horizontal distance in meters and p(x) is the vertical height in meters. What is p(2)?

OpenStudy (anonymous):

real equal real, imaginary equal imaginary \[3-x=2\] \[-2-y=-3\]

jimthompson5910 (jim_thompson5910):

(3-2i)-(x+yi) = 2-3i 3-2i-x-yi = 2-3i 3-x-2i-yi = 2-3i (3-x)+(-2i-yi) = 2-3i (3-x)+(-2-y)i = 2-3i Now equate the real parts to get 3-x = 2 -x = 2-3 -x = -1 x = 1 and equate the imaginary parts to get (-2-y)i = -3i -2-y = -3 -y = -3+2 -y = -1 y = 1